Evaluate using identities..
(99)3
Please tell me easy way to do it

Grade:12

1 Answers

Arun
5 years ago
Dear Priya
 
It is known that,
(a+b)3=a3+b3+3ab(a+b)and(a−b)3=a3−b3−3ab(a−b)

(i) (99)3 = (100 − 1)3
= (100)3 − (1)3 − 3(100) (1) (100 − 1)
= 1000000 − 1 − 300(99)
= 1000000 − 1 − 29700
= 970299